OpenStudy (anonymous):

The perimeter of a triangle is 76 cm. Side a of the triangle is twice as long as side b. Side c is 1 cm longer than side a. Find the length of each side.

OpenStudy (anonymous):

Do you have any ideas on how to go about solving this?

OpenStudy (anonymous):

Well, not really. I mean, I have some equations down, but I'm not sure if they are right. a = 2x b= x/2 c= x + 1

OpenStudy (anonymous):

A + B + C = P (Perimeter) Give me a minute, trying to remember triangles xD

OpenStudy (anonymous):

Okay! Thank you!

OpenStudy (anonymous):

A = 2b C = a + 1 This sound right?

OpenStudy (anonymous):

Yeah, that's what I have.

OpenStudy (anonymous):

So C = 2b + 1 2b + 1 = 76

OpenStudy (anonymous):

Try solving that.

OpenStudy (anonymous):

b = 37.5

OpenStudy (anonymous):

o.o wait.

OpenStudy (anonymous):

xD okie

OpenStudy (anonymous):

Okay so I tried it a different way. I kept the 37.5 and divided it by 2 again, because it's x/2. That gave me 18.75 and I added 1 to it. So I got 37.5, 18.75, and 19.75 which equals 76, but I'm not entirely sure that's 100% correct.

OpenStudy (anonymous):

Umm, looking at the problem again now, C is a + 1, not 2b + 1-- orz sorry

OpenStudy (anonymous):

I just substituted the A for 2b. ^^"

OpenStudy (anonymous):

oh oh omg sorry

OpenStudy (anonymous):

Aw, I gotta go. I have soccer practice. If you're still around, I'll try to help out later! >.<

OpenStudy (anonymous):

Oh, have fun haha! Thank you for helping! <3

OpenStudy (anonymous):

On my dad's phone atm, but I just got the answer! A is 30, B is 15, and C is 31 ^-^ Thanks for being so patient with me.

OpenStudy (anonymous):

Thank you again for taking the time to help me <3